Sketch a graph that has:

  1. 1 turning point, a maximum value and a y-intercept of (0,5). (3 pts)
  2. End behaviours of III → I, y-intercept of 2 and 2 x-intercepts. (3 pts)
  3. End behaviour of II → I, 1 x-intercept.
1
Expert's answer
2021-11-03T15:56:51-0400

1. A graph of parabola has 1 turning point. Let "f(x)=ax^2+bx+c."

If a function has a maximum value, then "a<0."

y-intercept: "f(0)=a(0)^2+b(0)+c=c"

Example;


"f(x)=-x^2+2x+5"



2. As "x\to-\infin, f(x)\to-\infin"

As "x\to\infin, f(x)\to\infin"

If "x_1, x_2" are x-intercepts then


"f(x)=a(x-x_1)(x-x_2)^2, a>0"

"f(0)=-ax_1(-x_2)^2=2"

Let "a=2, x_1=-1, x_2=1"


"f(x)=2(x+1)(x-1)^2"



3. As "x\to-\infin, f(x)\to\infin"

As "x\to\infin, f(x)\to\infin"

If "x_1" is x-intercepts then


"f(x)=a(x-x_1)^2, a>0"

Let "a=1, x=1"


"f(x)=(x-1)^2"